# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 1.1 Motivation
Humans are the foremost actors in various social activities, and it is imperative to equip AI with an understanding of humans for contributing to society. Consequently, there are many human-centric tasks positioning the epicenter of research. Among them, 3D Human Pose Estimation (HPE) and Human Mesh Recovery (HMR) represent crucial tasks in the field of computer vision to interpret the status and behavior of humans in complex real-world environments. 3D human pose estimation can accurately predict human body keypoint coordinates in three-dimensional space. This approach provides more comprehensive and accurate spatial information when compared to its 2D counterpart, thereby facilitating a better understanding of complex human behaviors in higher-level computer vision applications duan2022revisiting ; zhang2022voxeltrack ; zhu2022multilevel . Human mesh recovery reconstructs a three-dimensional digital model of the body, which captures details such as shape you2023co ; tripathi20233d , gestures fan2023hold , clothing dai2023cloth2body ; tang2023high , and facial expressions feng2023learning , thus offering direct insights into human interactions with the physical world. 3D pose estimation and mesh recovery have a broad range of applications, such as security and surveillance wang2021horeid , human-computer interaction liu2022arhpe , autonomous driving zheng2022multi ; wang2023learning , and virtual reality weng2019photo . With advanced deep learning technology, 3D human pose estimation and mesh recovery With advanced deep learning technology, 3D human pose estimation and mesh recovery have garnered increasing attention in recent years. 3D pose estimation has evolved from concentrating on single individuals to encompassing multiple persons with more varied data inputs. In human mesh recovery, advancements have been made in terms of data inputs and in capturing more intricate details. The augmentation in explicit model parameters allows for a more nuanced representation of the human body, and the expansion in parameter types facilitates the portrayal of finer surfaces. With the progress of implicit rendering and its incorporation into human mesh recovery, more flexible body representations are achieved. However, both 3D pose estimation and mesh recovery face significant challenges, such as multi-person scenarios, self-occlusion issues, and the detailed reconstruction of bodies. Thus, a systematic and comprehensive review of recent advancements in 3D human pose estimation and mesh recovery is essential.
@@ 1.2 Scope of this survey
Over the past five years, numerous reviews have been on 3D human pose estimation and mesh recovery. The reviews liu2022recent ; zheng2023deep primarily concentrate on 2D and 3D pose estimation, yet they devote less attention to HMR-related literature. The survey tian2023recovering provides a thorough review exclusively dedicated to mesh recovery, focusing on methods based on explicit models. The survey chen2021towards is distinctly centered on mesh recovery using implicit rendering techniques; however, it falls short in offering an exhaustive overview of the latest implicit rendering methods and systems for 3D pose estimation and mesh recovery. These reviews only marginally explore 3D pose estimation and mesh recovery technologies but scarcely delve into their applications in other computer vision tasks and future challenges.
@@
The main contributions of this work compared to the existing literature can be summarized as follows: (1) To the best of our knowledge, this survey is arguably the first to comprehensively cover deep learning methods for 3D human pose estimation, including both single-person and multi-person approaches, as well as human mesh recovery, encompassing methods based on explicit models and implicit representations. (2) Unlike existing reviews, we have not overlooked the role of implicit representations in the methods, particularly with the recent rapid advances in implicit rendering. This approach can produce detailed outputs, including clothed human figures with expressions, movements, and other intricacies essential for achieving photorealism. (3) This paper comprehensively reviews the most recent developments in deep learning for 3D pose estimation and mesh recovery, providing readers with a detailed overview of cutting-edge methodologies. Additionally, it explores how these advancements contribute to various other computer vision tasks and delves into the challenges within this domain. The structure of this paper is as follows: Section 2 introduces the sensor type and representation for the human body that have been widely used. Section 3 presents the overview of deep learning for 3D human pose estimation and mesh recovery. Section 4 surveys existing single person and multi-person 3D pose estimation methods. Section 5 reviews the methods for human mesh recovery, including template-based and template-free methods. Section 6 introduces the evaluation metrics and datasets for the respective tasks. Moreover, Section 7 discusses the applications, along with their impact on other computer vision tasks. Finally, Section 8 concludes the paper. Additionally, we host a regularly updated project page, which can be accessed at: https://github.com/liuyangme/SOTA-3DHPE-HMR.
@@ 2.1.1 Active sensors
Active sensors operate by emitting a set of signals and measuring by detecting their reflections, such as Motion Capture (MoCap) systems with reflective markers einfalt2023uplift , tactile sensing luo2021intelligent , Time of Flight (ToF) cameras ruget2022pixels2pose ; pandey2019volumetric , and Radio Frequency (RF) technologies ren2022gopose ; li2022unsupervised . MoCap and tactile devices are only suitable for cooperative targets. Active cameras generally cannot be used outdoors, and using multiple active devices simultaneously may lead to mutual interference. Thus, these devices have limited application scenarios.
@@ 2.1.2 Passive sensors
Passive sensors do not actively emit any signals during the measurement; instead, they rely on signals from the objects or natural sources, including Inertial Measurement Units (IMUs) ponton2023sparseposer ; huang2020deepfuse and image sensors zou2022human ; xu2020eventcap . Among these, RGB image sensors are particularly notable for being simple, user-friendly, adaptable to various environments, and capable of capturing high-resolution color images. Additionally, multiple RGB image sensors can be combined to form multi-view systems jiang2023probabilistic ; shuai2022adaptive ; huang2021dynamic . In this survey, considering the widespread applicability of RGB image sensors and the length limitations of the article, we focus on 3D human pose estimation and mesh recovery using RGB image sensors.
@@ 2.2 Representation for human body
3D pose estimators and mesh reconstructors can generate corresponding outcomes from the sensor input described previously. The 3D human pose estimation output comprises 3D coordinates detailing positions and joint orientations of the human body, representing the spatial location of each joint (e.g., head, neck, shoulders, elbows, knees), and the skeleton maps the interconnections between joints. Typically, these coordinates are expressed in a global coordinate system or relative to the camera’s coordinate system. A keypoints tree structure is commonly employed to illustrate the human pose, where the tree nodes represent the joints and the edges denote the connections between joints. However, the skeleton-based human pose representation method does not provide detailed information about the body’s surface. The output of human mesh recovery is generally a 3D body model, which encompasses a comprehensive depiction of the body’s shape and surface details and offers a more enriched representation. Statistical-based models are widely used in human mesh representations, such as SCAPE anguelov2005scape and SMPL loper2015smpl . SMPL is a learnable skin-vertex model representing the human body as a 3D mesh with a topological structure. The pose and shape of the body are described by pose parameters $\theta$ and shape parameters $\beta$. The pose parameters control the joint angles and global posture, and the shape parameters determine the body’s shape. There are several models based on SMPL to expand the representational capabilities, such as the MANO model (SMPL+H) romero2022embodied with hand representation and the FLAME model li2017learning with facial representation. SMPL-X pavlakos2019expressive is a comprehensive model that simultaneously captures the human body, face, and hands by incorporating the FLAME head model li2017learning and the MANO hand model romero2022embodied . H4D jiang2022h4d can represent the dynamic human shape and pose by building upon the prior knowledge from the SMPL model and extending its capabilities by incorporating a temporal dimension. In addition to the explicit model representations mentioned above, recent years have seen the development of methods based on implicit models, which offer a more flexible representation of the human body through non-parametric models. For example, to address the limitations of previous representation models, Zheng et al. zheng2021pamir designed the Parametric Model-Conditioned Implicit Representation (PaMIR), which utilizes the 2D image feature map and corresponding SMPL feature volume to generate an implicit surface representation.
@@ 3 Overview of Deep Learning for 3D HPE and HMR
This section provides an overview of the deep learning framework for 3D human pose estimation and mesh recovery. As shown in Fig.2, there are four components in deep learning-based systems. First, data are collected from various sensors, including RGB images, depth images, RF signals, IMUs, and so on. Second, the input data are processed through a deep learning model, which typically consists of an encoder and a decoder. The encoder extracts representational features from the input data, such as those obtained using architectures like ResNet he2016deep or HRNet wang2020deep . The decoder, which could be based on frameworks like MLP (Multi-Layer Perceptron) tolstikhin2021mlp or Transformer dosovitskiy2020image , then outputs the body’s 3D pose and reconstruction model. This structure enables the model to efficiently process and translate complex input data into detailed and accurate 3D representations of human pose and mesh. Thirdly, various learning methods can be selected during the model learning process. In addition to fully supervised learning, to alleviate data dependency approaches such as weakly supervised learning yang2023camerapose ; zanfir2020weakly , unsupervised learning chai2023global ; yu2021skeleton2mesh , and few-shot learning mu2023actorsnerf ; benzine2020pandanet are employed. To reduce the model size, techniques such as knowledge distillation yang2023effective ; tripathi2020posenet3d , model pruning liu2019effective , and parameter quantization choi2021mobilehumanpose can be applied. Furthermore, methodologies such as meta-learning cho2021camera and reinforcement learning gong2022posetriplet can also be incorporated, allowing the model to adapt to different scenarios and data constraints. Finally, the deep learning model outputs the results of 3D human pose estimation and mesh recovery. These results can be represented in various forms, including keypoints hassan2023regular ; zhao2023poseformerv2 , mesh loper2015smpl ; romero2022embodied ; cai2023smpler , and voxels li2020monocular ; onizuka2020tetratsdf ; tang2023high . These representations contribute to a comprehensive understanding of the human body in three dimensions. In the following sections, we will delve into the specifics of 3D human pose estimation and mesh recovery, categorizing and elaborating on each aspect.
@@
Single Person in Images Solving Depth Ambiguity $\bullet$ Optical-aware: VI-HC wei2019view , Ray3D zhan2022ray3d
$\bullet$ Appropriate feature representation: HEMlets zhou2021hemlets
@@
Solving Body Structure Understanding $\bullet$ Joint aware: JRAN zheng2020joint
$\bullet$ Limb aware: Wu et al. wu2021limb , Deep grammar network xu2021monocular
$\bullet$ Orientation keypoints: Fisch et al. fisch2021orientation
$\bullet$ Graph-based: Liu et al. liu2019feature , LCN ci2020locally , Modulated-GCN zou2021modulated , Skeletal-GNN zeng2021learning , HopFIR zhai2023hopfir , RS-Net hassan2023regular
Solving Occlusion Problems $\bullet$ Learnable-triangulation iskakov2019learnable
$\bullet$ RPSM qiu2019cross
$\bullet$ Lightweight multi-view remelli2020lightweight
$\bullet$ AdaFuse zhang2021adafuse
$\bullet$ Bartol et al. bartol2022generalizable
$\bullet$ 3D pose consensus luvizon2022consensus
$\bullet$ Probabilistic triangulation module jiang2023probabilistic
@@
Solving Data Lacking $\bullet$ Unsupervised learning: Kudo et al. kudo2018unsupervised , Chen et al. chen2019unsupervised , ElePose wandt2022elepose
$\bullet$ Self-supervised learning: EpipolarPose kocabas2019self , Wang et al. wang20193d , MRP-Net kundu2022uncertainty , PoseTriplet gong2022posetriplet
$\bullet$ Weakly-supervised learning: Hua et al. hua2022weakly , CameraPose yang2023camerapose
$\bullet$ Transfer learning: Adaptpose gholami2022adaptpose
in Videos Solving Single-frame Limitation $\bullet$ VideoPose3D pavllo20193d
$\bullet$ PoseFormer zheng20213d
$\bullet$ UniPose+ artacho2021uniposeplus
$\bullet$ MHFormer li2022mhformer
$\bullet$ MixSTE zhang2022mixste
$\bullet$ Honari et al. honari2022temporal
$\bullet$ HSTFormer qian2023hstformer
$\bullet$ STCFormer tang20233d
@@
Solving Real-time Problems $\bullet$ Temporally sparse sampling: Einfalt et al. einfalt2023uplift
$\bullet$ Spatio-temporal sparse sampling: MixSynthFormer sun2023mixsynthformer
@@
Solving Body Structure Understanding $\bullet$ Motion loss: Wang et al. wang2020motion
$\bullet$ Human-joint affinity: DG-Net zhang2021learning
$\bullet$ Anatomy-aware: Chen et al. chen2021anatomy
$\bullet$ Part aware attention: Xue et al. xue2022boosting
@@
Solving Occlusion Problems $\bullet$ Optical-flow consistency constraint: Cheng et al. cheng2019occlusion
$\bullet$ Multi-view: MTF-Transformer shuai2022adaptive
Solving Data Lacking $\bullet$ Unsupervised learning: Yu et al. yu2021towards
$\bullet$ Weakly-supervised learning: Chen et al. chen2019weakly
$\bullet$ Semi-supervised learning: MCSS mitra2020multiview
$\bullet$ Self-supervised learning: Kundu et al. kundu2020self , P-STMO shan2022p
$\bullet$ Meta-learning: Cho et al. cho2021camera
$\bullet$ Data augmentation: PoseAug gong2021poseaug , Zhang et al. zhang2023learning
@@
Multi-person Top-down Solving Real-time Problems $\bullet$ Multi-view: Chen et al. chen2020cross
$\bullet$ Whole body: AlphaPose fang2022alphapose
Solving Representation Limitation $\bullet$ VoxelTrack zhang2022voxeltrack
Solving Occlusion Problems $\bullet$ Wu et al. wu2021graph
Solving Data Lacking $\bullet$ Single-shot: PandaNet benzine2020pandanet
$\bullet$ Optical-aware: Moon et al. moon2019camera
Bottom-up Solving Real-time Problems $\bullet$ Fabbri et al. fabbri2020compressed
Solving Supervisory Limitation $\bullet$ HMOR wang2020hmor
@@
Solving Data Lacking $\bullet$ Single-shot: SMAP zhen2020smap , Benzine et al. benzine2021single
Solving Occlusion Problems $\bullet$ Mehta et al. mehta2018single
$\bullet$ LCR-Net++ rogez2019lcr
Others Single Stage $\bullet$ Jin et al. jin2022single
Top-down + Bottom-up $\bullet$ Cheng et al. cheng2022dual

### v2

@@ 1.1 Motivation
Humans are the foremost actors in various social activities, and it is imperative to equip AI with an understanding of humans for contributing to society. Consequently, there are many human-centric tasks positioning the epicenter of research. Among them, 3D Human Pose Estimation (HPE) and Human Mesh Recovery (HMR) represent crucial tasks in the field of computer vision to interpret the status and behavior of humans in complex real-world environments.
3D human pose estimation can accurately predict human body keypoint coordinates in three-dimensional space. This approach provides more comprehensive and accurate spatial information when compared to its 2D counterpart, thereby facilitating a better understanding of complex human behaviors in higher-level computer vision applications [1, 2, 3, 4]. Human mesh recovery reconstructs a three-dimensional digital model of the body, which captures details such as shape [5, 6], gestures [7], clothing [8, 9], and facial expressions [10], thus offering direct insights into human interactions with the physical world. 3D pose estimation and mesh recovery have a broad range of applications, such as security and surveillance [11], human-computer interaction [12, 13], autonomous driving [14, 15], and virtual reality [16].
With advanced deep learning technology, 3D human pose estimation and mesh recovery have garnered increasing attention in recent years. 3D pose estimation has evolved from concentrating on single individuals to encompassing multiple persons with more varied data inputs. In human mesh recovery, advancements have been made in terms of data inputs and in capturing more intricate details. The augmentation in explicit model parameters allows for a more nuanced representation of the human body, and the expansion in parameter types facilitates the portrayal of finer surfaces. With the progress of implicit rendering and its incorporation into human mesh recovery, more flexible body representations are achieved. However, both 3D pose estimation and mesh recovery face significant challenges, such as multi-person scenarios, self-occlusion issues, and the detailed reconstruction of bodies. Thus, a systematic and comprehensive review of recent advancements in 3D human pose estimation and mesh recovery is essential.
@@ 1.2 Scope of this survey
Over the past five years, numerous reviews have been on 3D human pose estimation and mesh recovery. The reviews [17, 18] primarily concentrate on 2D and 3D pose estimation, yet they devote less attention to HMR-related literature. The survey [19] provides a thorough review exclusively dedicated to mesh recovery, focusing on methods based on explicit models. The survey [20] is distinctly centered on mesh recovery using implicit rendering techniques; however, it falls short in offering an exhaustive overview of the latest implicit rendering methods and systems for 3D pose estimation and mesh recovery. These reviews only marginally explore 3D pose estimation and mesh recovery technologies but scarcely delve into their applications in other computer vision tasks and future challenges.
@@
The main contributions of this work compared to the existing literature can be summarized as follows:
(1) To the best of our knowledge, this survey is arguably the first to comprehensively cover deep learning methods for 3D human pose estimation, including both single-person and multi-person approaches, as well as human mesh recovery, encompassing methods based on explicit models and implicit representations.
(2) Unlike existing reviews, we have not overlooked the role of implicit representations in the methods, particularly with the recent rapid advances in implicit rendering. This approach can produce detailed outputs, including clothed human figures with expressions, movements, and other intricacies essential for achieving photorealism.
(3) This paper comprehensively reviews the most recent developments in deep learning for 3D pose estimation and mesh recovery, providing readers with a detailed overview of cutting-edge methodologies. Additionally, it explores how these advancements contribute to various other computer vision tasks and delves into the challenges within this domain.
The structure of this paper is as follows: Section 2 introduces the sensor type and representation for the human body that have been widely used. Section 3 presents the overview of deep learning for 3D human pose estimation and mesh recovery. Section 4 surveys existing single person and multi-person 3D pose estimation methods. Section 5 reviews the methods for human mesh recovery, including template-based and template-free methods. Section 6 introduces the evaluation metrics and datasets for the respective tasks. Moreover, Section 7 discusses the applications, along with their impact on other computer vision tasks. Finally, Section 8 concludes the paper. Fig.2 shows the taxonomy of deep learning methods for 3D HPE and HMR. Additionally, we host a regularly updated project page, which can be accessed at: https://github.com/liuyangme/SOTA-3DHPE-HMR.
Figure 2: A taxonomy of deep learning methods for 3D HPE and HMR.
@@ 2.1.1 Active sensors
Active sensors operate by emitting a set of signals and measuring by detecting their reflections, such as Motion Capture (MoCap) systems with reflective markers [21], tactile sensing [22], Time of Flight (ToF) cameras [23, 24], and Radio Frequency (RF) technologies [25, 26]. MoCap and tactile devices are only suitable for cooperative targets. Active cameras generally cannot be used outdoors, and using multiple active devices simultaneously may lead to mutual interference. Thus, these devices have limited application scenarios.
@@ 2.1.2 Passive sensors
Passive sensors do not actively emit any signals during the measurement; instead, they rely on signals from the objects or natural sources, including Inertial Measurement Units (IMUs) [27, 28] and image sensors [29, 30]. Among these, RGB image sensors are particularly notable for being simple, user-friendly, adaptable to various environments, and capable of capturing high-resolution color images. Additionally, multiple RGB image sensors can be combined to form multi-view systems [31, 32, 33].
In this survey, considering the widespread applicability of RGB image sensors and the length limitations of the article, we focus on 3D human pose estimation and mesh recovery using RGB image sensors.
@@ 2.2 Representation for human body
3D pose estimators and mesh reconstructors can generate corresponding outcomes from the sensor input described previously. The 3D human pose estimation output comprises 3D coordinates detailing positions and joint orientations of the human body, representing the spatial location of each joint (e.g., head, neck, shoulders, elbows, knees), and the skeleton maps the interconnections between joints. Typically, these coordinates are expressed in a global coordinate system or relative to the camera’s coordinate system. A keypoints tree structure is commonly employed to illustrate the human pose, where the tree nodes represent the joints and the edges denote the connections between joints.
On the other hand, the skeleton-based human pose representation method does not provide detailed information about the body’s surface. The output of human mesh recovery is generally a 3D body model, which encompasses a comprehensive depiction of the body’s shape and surface details and offers a more enriched representation. Statistical-based models are widely used in human mesh representations, such as SCAPE [34] and SMPL [35]. SMPL is a learnable skin-vertex model representing the human body as a 3D mesh with a topological structure. The pose and shape of the body are described by pose parameters $\theta$ and shape parameters $\beta$. The pose parameters control the joint angles and global posture, and the shape parameters determine the body’s shape. There are several models based on SMPL to expand the representational capabilities, such as the MANO model (SMPL+H) [36] with hand representation and the FLAME model [37] with facial representation. SMPL-X [38] is a comprehensive model that simultaneously captures the human body, face, and hands by incorporating the FLAME head model [37] and the MANO hand model [36]. H4D [39] can represent the dynamic human shape and pose by building upon the prior knowledge from the SMPL model and extending its capabilities by incorporating a temporal dimension.
In addition to the explicit model representations mentioned above, recent years have seen the development of methods based on implicit models, which offer a more flexible representation of the human body through non-parametric models. The voxel-based model is suitable for volume rendering and physical simulation. However, its resolution is limited by the size of the voxels, which may result in larger storage requirements. Varol et al. [40] proposed an end-to-end network architecture, BodyNet, capable of predicting the volumetric shape of the human body from a single image. Implicit reconstruction transforms inputs into implicit function representations, facilitating the reconstruction of high-quality three-dimensional mesh models from irregular and noisy data. To address the ill-posed problem of reconstructing a 3D mesh from images, Onizuka et al. proposed the TetraTSDF (Tetrahedral Truncated Signed Distance Function) model [41]. TetraTSDF is a tetrahedral representation model capable of accurately recovering intricate human body shapes, even when the subject is wearing loose clothing. Zheng et al. [42] designed the Parametric Model-Conditioned Implicit Representation (PaMIR), which utilizes the 2D image feature map and corresponding SMPL feature volume to generate an implicit surface representation.
@@ 3 Overview of Deep Learning for 3D HPE and HMR
This section provides an overview of the deep learning framework for 3D human pose estimation and mesh recovery. As shown in Fig.3, there are four components in deep learning-based systems. First, data are collected from various sensors, including RGB images, depth images, RF signals, IMUs, and so on. Second, the input data are processed through a deep learning model, which typically consists of an encoder and a decoder. The encoder extracts representational features from the input data, such as those obtained using architectures like ResNet [43] or HRNet [44]. The decoder, which could be based on frameworks like MLP (Multi-Layer Perceptron) [45] or Transformer [46], then outputs the body’s 3D pose and reconstruction model. This structure enables the model to efficiently process and translate complex input data into detailed and accurate 3D representations of human pose and mesh. Thirdly, various learning methods can be selected during the model learning process. In addition to fully supervised learning, to alleviate data dependency approaches such as weakly supervised learning [47, 48], unsupervised learning [49, 50], and few-shot learning [51, 52] are employed. To reduce the model size, techniques such as knowledge distillation [53, 54], model pruning [55], and parameter quantization [56] can be applied. Furthermore, methodologies such as meta-learning [57] and reinforcement learning [58] can also be incorporated, allowing the model to adapt to different scenarios and data constraints. Finally, the deep learning model outputs the results of 3D human pose estimation and mesh recovery. These results can be represented in various forms, including keypoints [59, 60], mesh [35, 36, 61], and voxels [62, 41, 9]. These representations contribute to a comprehensive understanding of the human body in three dimensions. In the following sections, we will delve into the specifics of 3D human pose estimation and mesh recovery, categorizing and elaborating on each aspect.
@@
Single Person in Images Solving Depth Ambiguity $\bullet$ Optical-aware: VI-HC [63], Ray3D [64]
$\bullet$ Appropriate feature representation: HEMlets [65]
@@
Solving Body Structure Understanding $\bullet$ Joint aware: JRAN [66]
$\bullet$ Limb aware: Wu et al. [67], Deep grammar network [68]
$\bullet$ Orientation keypoints: Fisch et al. [69]
$\bullet$ Graph-based: Liu et al. [70], LCN [71], Modulated-GCN [72], Skeletal-GNN [73], HopFIR [74], RS-Net [59]
Solving Occlusion Problems $\bullet$ Learnable-triangulation [75]
$\bullet$ RPSM [76]
$\bullet$ Lightweight multi-view [77]
$\bullet$ AdaFuse [78]
$\bullet$ Bartol et al. [79]
$\bullet$ 3D pose consensus [80]
$\bullet$ Probabilistic triangulation module [31]
@@
Solving Data Lacking $\bullet$ Unsupervised learning: Kudo et al. [81], Chen et al. [82], ElePose [83]
$\bullet$ Self-supervised learning: EpipolarPose [84], Wang et al. [85], MRP-Net [86], PoseTriplet [58]
$\bullet$ Weakly-supervised learning: Hua et al. [87], CameraPose [47]
$\bullet$ Transfer learning: Adaptpose [88]
in Videos Solving Single-frame Limitation $\bullet$ VideoPose3D [89]
$\bullet$ PoseFormer [90]
$\bullet$ UniPose+ [91]
$\bullet$ MHFormer [92]
$\bullet$ MixSTE [93]
$\bullet$ Honari et al. [94]
$\bullet$ HSTFormer [95]
$\bullet$ STCFormer [96]
@@
Solving Real-time Problems $\bullet$ Temporally sparse sampling: Einfalt et al. [21]
$\bullet$ Spatio-temporal sparse sampling: MixSynthFormer [97]
@@
Solving Body Structure Understanding $\bullet$ Motion loss: Wang et al. [98]
$\bullet$ Human-joint affinity: DG-Net [99]
$\bullet$ Anatomy-aware: Chen et al. [100]
$\bullet$ Part aware attention: Xue et al. [101]
@@
Solving Occlusion Problems $\bullet$ Optical-flow consistency constraint: Cheng et al. [102]
$\bullet$ Multi-view: MTF-Transformer [32]
Solving Data Lacking $\bullet$ Unsupervised learning: Yu et al. [103]
$\bullet$ Weakly-supervised learning: Chen et al. [104]
$\bullet$ Semi-supervised learning: MCSS [105]
$\bullet$ Self-supervised learning: Kundu et al. [106], P-STMO [107]
$\bullet$ Meta-learning: Cho et al. [57]
$\bullet$ Data augmentation: PoseAug [108], Zhang et al. [109]
@@
Multi-person Top-down Solving Real-time Problems $\bullet$ Multi-view: Chen et al. [110]
$\bullet$ Whole body: AlphaPose [111]
Solving Representation Limitation $\bullet$ VoxelTrack [2]
Solving Occlusion Problems $\bullet$ Wu et al. [112]
Solving Data Lacking $\bullet$ Single-shot: PandaNet [52]
$\bullet$ Optical-aware: Moon et al. [113]
Bottom-up Solving Real-time Problems $\bullet$ Fabbri et al. [114]
Solving Supervisory Limitation $\bullet$ HMOR [115]
@@
Solving Data Lacking $\bullet$ Single-shot: SMAP [116], Benzine et al. [117]
Solving Occlusion Problems $\bullet$ Mehta et al. [118]
$\bullet$ LCR-Net++ [119]
Others Single Stage $\bullet$ Jin et al. [120]
Top-down + Bottom-up $\bullet$ Cheng et al. [121]